OpenStudy (anonymous):

Identify the conclusion of the following conditional: A number is divisible by 3 if the sum of the digits of the number is divisible by 3. The number is odd. If the sum of the digits of a number is divisible by 3, then the number is divisible by 3. The number is divisible by 3. The sum of the digits of the number is divisible by 3.

OpenStudy (anonymous):

that was wrong rewrite as "if the sum of the digits of the number is divisible by 3. then the number is divisible by 3" the conclusion is the part that comes after the word "then"

OpenStudy (jiteshmeghwal9):

A number is divisible by 3 if the sum of the digits of a number is divisible by 3. Example:- 27=7+2=9 which is divisible by 3.

OpenStudy (anonymous):

who is this insane

OpenStudy (jusaquikie):

12= 1+2=3 Divisible by 3; 99 = 9+9=18 = 1+8=9 divisible by 3; 1234= 1+2+3+4=10 = 1+0=1 not divisible by 3

OpenStudy (jusaquikie):

nine works the same way FYI

OpenStudy (anonymous):

Determine whether the conditional and its converse are both true. If both are true, combine them as a biconditional. If either is false, provide a counterexample. If an angle is a right angle, its measure is 90. If an angle measure is 90, the angle is a right angle. One statement is false. If an angle measure is 90, the angle may be an obtuse angle. Both statements are true. An angle is a right angle if and only if its measure is 90. One statement is false. If an angle is a right angle, its measure may be 180. Both statements are true. The measure of an angle is 90 if and only if it is not a right angle

OpenStudy (anonymous):

moderators please stop it or try to do something

OpenStudy (anonymous):

@mattfeury please sir check this out he/she is annoying too much

OpenStudy (mattfeury):

we got it. just click Report Abuse when you see this, rather than replying off-topic in someone's questions. Thanks :)

OpenStudy (anonymous):

i have already done that now i am unable to report

OpenStudy (anonymous):

Read the following two statements. Then, if possible, use the Law of Detachment to draw a conclusion. If two figures are congruent, their areas are equal. The area of ABCD equals the area of PQRS. not possible Figure ABCD is congruent to figure PQRS. Figure ABCD is not congruent to figure PQRS. If the areas of two figures are equal, the figures are congruent